# An extension of Priestley duality to fuzzy topologies and positive MV-algebras

### v1

@@ 4 The extension of Priestley Duality
The restrictions of $\operatorname{Clop}_{\uparrow}$ and $\Upsilon$ form a duality between $\mathcal{MV}^{\operatorname{lcc}}_{H}$ and $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}^{\operatorname{lcc}}$.
5 Comparing the extensions of Priestley and Stone dualities
In this section, we show that the duality developed in the previous section extends the Stone duality for MV-algebras, similar to what occurs between the corresponding dualities in the classical case.
Let $A$ be an MV-lattice and $B\subseteq[0,1]^{H_{A}}$ the MV-algebra generated by $\tilde{A}$. Since $B$ is a semisimple MV-algebra, $B$ is isomorphic to an MV-subalgebra $\bar{B}$ of the MV-algebra of fuzzy subsets $0,1]^{MaxB}$ and $\bar{B}$ is a base of an MV-topology $\Omega_{B}$ over $\operatorname{Max}B$. In this way, we obtain the Stone MV-space $(\operatorname{Max}B,\Omega_{B})$, i.e., the image of $B$ under the functor $\operatorname{Max}:\mathcal{MV}\longrightarrow{}^{\operatorname{MV}}\!\mathcal{T}\!\!\operatorname{op}$ defined in [15].

### v2

@@ 4 The extension of Priestley Duality
The restrictions of $\operatorname{Clop}_{\uparrow}$ and $\Upsilon$ form a duality between $\mathcal{MV}^{\operatorname{lcc}}_{+H}$ and $^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}^{\operatorname{lcc}}$.
5 Comparing the dualities
In this section, we show that the duality developed in the previous section extends both classical Priestley Duality and the Stone-type duality for MV-algebras.
On the one hand, our duality is a proper extension of the classical one, in the sense that the categories involved in the latter are full subcategories of the ones involved in our duality, and the restrictions of the functors defined in the previous section to such subcategories yield the classical Priestley Duality.
On the other hand, just like classical Priestley Duality extends Stone Duality, our duality extends the one between lcc MV-algebras and Stone MV-spaces of [14].
In order to illustrate the relationship between $(\Upsilon,\operatorname{Clop}_{\uparrow})$ and classical Priestley Duality, we first observe that, if $A$ is a bounded distributive lattice thought of as an MV-lattice, we have that ++ and $\cdot$ coincide with $\vee$ and $\wedge$ respectively. Therefore, if $f:A\to[0,1]$ is an MV-lattice homomorphism, for all $a\in A$, $f(a)$ must be an additively idempotent element of $[0,1]$, and this implies that $f[A]=\{0,1\}$. So, homomorphisms of $A$ with values in $[0,1]$ are in bijective correspondence with those with values in $\{0,1\}$, and it is immediate to see that everything collapses to classical Priestley Duality.
For what concerns the relationship of our duality with the one in [14], let $A$ be an MV-lattice and $B\subseteq[0,1]^{H_{A}}$ the MV-algebra generated by $\tilde{A}$. Since $B$ is a semisimple MV-algebra, $B$ is isomorphic to an MV-subalgebra $\bar{B}$ of the MV-algebra of fuzzy subsets $0,1]^{MaxB}$ and $\bar{B}$ is a base of an MV-topology $\Omega_{B}$ over $\operatorname{Max}B$. In this way, we obtain the Stone MV-space $(\operatorname{Max}B,\Omega_{B})$, i.e., the image of $B$ under the functor $\operatorname{Max}:\mathcal{MV}\longrightarrow{}^{\operatorname{MV}}\!\mathcal{T}\!\!\operatorname{op}$ defined in [14].
It is worth noticing that the fuzzy version of Stone duality presented in [14] was the starting point of a sheaf representation of a large class of MV-algebra and a sheaf embedding of all MV-algebras, published in [7]. The introduction of positive MV-algebra is very recent, therefore, the algebraic theory is not sufficiently developed yet to allow a similar sheaf representation as an easy consequence of our result. However, the good behaviour of the structure suggests that such a result may be possible in the next future.
@@ 5 Comparing the dualities
Summing up, we have the following diagram of categorical full embeddings and dualities, with obvious meaning of the arrows, $\mathcal{BDL}$ being the category of bounded distributive lattices with their homomorphisms.
$\mathcal{P}\!\!\operatorname{ries}$$\mathcal{BDL}$$^{\operatorname{MV}}\!\mathcal{P}\!\!\operatorname{ries}^{\operatorname{lcc}}$$\mathcal{MV}^{\operatorname{lcc}}_{+H}$$\mathcal{S}\!\operatorname{tone}$$\mathcal{B}\!\operatorname{oole}$$^{\operatorname{MV}}\!\!\mathcal{S}\!\operatorname{tone}$$\mathcal{MV}^{\operatorname{lcc}}$}